OpenStudy (anonymous):

(3-square root of 7)(3+square root of 7)

OpenStudy (anonymous):

apply a^2 -b^2 formula.. i.e (a+b)(a-b)=a^2-b^2 (3)^2-(root 7)^2=9-7 =2;; cheers!!!